determine the price and quantity if Q=200-2P+200=-50+3P


1
Expert's answer
2021-08-24T14:07:57-0400
"Q_D=200-2p"

"Q_S=-50+3p"

"200-2p=-50+3p"

"5p=250"

"p=50"

"Q=100"